# Gauge coupling unification and proton decay via 45 Higgs boson in SU(5) GUT

### v2

@@ 2 Model: $SU(5)$ GUT with a 45 Higgs Field
[…]
@@ 3 Gauge Coupling Unification
$Q_{i}=\begin{pmatrix}u_{i}\\ d_{i}\end{pmatrix}_{L}$ $(3,2)_{1/6}$ $2$ $3$ $1/5$
$u_{i}^{~{}c}=(u_{i}^{~{}c})_{L}$ $(\overline{3},1)_{-2/3}$ $1$ 00 $8/5$
$d_{i}^{~{}c}=(d_{i}^{~{}c})_{L}$ $(\overline{3},1)_{1/3}$ $1$ 00 $2/5$
$L_{i}=\begin{pmatrix}\nu_{ei}\\ e_{i}\end{pmatrix}_{L}$ $(1,2)_{-1/2}$ 00 $1$ $3/5$
$e_{i}^{~{}c}=(e_{i}^{~{}c})_{L}$ $(1,1)_{1}$ 00 00 $6/5$
$H=\begin{pmatrix}\phi_{+}\\ \phi_{0}\end{pmatrix}$ $(1,2)_{-1/2}$ 00 $1/6$ $1/10$
$H^{c}=\begin{pmatrix}H_{1}\\ H_{2}\\ H_{3}\end{pmatrix}$ $(\overline{3},1)_{1/3}$ $1/6$ 00 $1/15$
$(X^{\mu},Y^{\mu})$ $(3,2)_{-5/6}$ $-7$ $-21/2$ $-35/2$
$\Sigma_{8}$ $(8,1)_{0}$ $1/2$ 00 00
$\Sigma_{3}$ $(1,3)_{0}$ 00 $1/3$ 00
$\Sigma_{24}$ $(1,1)_{0}$ 00 00 00
$(1,2)_{-1/2}$ $(1,2)_{-1/2}$ 00 $1/6$ $1/10$
$(\overline{3},1)_{1/3}$ $(\overline{3},1)_{1/3}$ $1/6$ 00 $1/15$
$(\overline{3},3)_{1/3}$ $(\overline{3},3)_{1/3}$ $1/2$ $2$ $1/5$
$(3,1)_{-4/3}$ $(3,1)_{-4/3}$ $1/6$ 00 $16/15$
@@
$\alpha_{1}^{-1}(M_{Z})=\alpha_{5}^{-1}(\Lambda)-\frac{1}{2\pi}\biggl{[}$ $\frac{41}{10}\log\frac{M_{Z}}{\Lambda}+\frac{1}{15}\log\frac{M_{H_{C}}}{\Lambda}-\frac{35}{2}\log\frac{M_{XY}}{\Lambda}+\frac{1}{10}\log\frac{M(1,2)}{\Lambda}$
@@
$\alpha_{2}^{-1}(M_{Z})-\frac{2}{12\pi}=\alpha_{5}^{-1}(\Lambda)-\frac{1}{2\pi}\biggl{[}$ $-\frac{19}{6}\log\frac{M_{Z}}{\Lambda}-\frac{21}{2}\log\frac{M_{XY}}{\Lambda}+\frac{1}{3}\log\frac{M_{\Sigma}}{\Lambda}+\frac{1}{6}\log\frac{M(1,2)}{\Lambda}$
@@
$\alpha_{3}^{-1}(M_{Z})-\frac{3}{12\pi}=\alpha_{5}^{-1}(\Lambda)-\frac{1}{2\pi}\biggl{[}$ $-7\log\frac{M_{Z}}{\Lambda}+\frac{1}{6}\log\frac{M_{H_{C}}}{\Lambda}-7\log\frac{M_{XY}}{\Lambda}+\frac{1}{2}\log\frac{M_{\Sigma}}{\Lambda}$
@@
where $M(a,b)$ denotes the mass of the 45 Higgs field component in $(a,b)$-representation of SM SU(3)$\times$×SU(2). From these equations, we derive the following two independent equations:
@@
$-5\alpha_{1}^{-1}(M_{Z})+3\alpha_{2}^{-1}(M_{Z})+2\alpha_{3}^{-1}(M_{Z})=-\frac{1}{2\pi}\biggl{[}$ $-42\log\frac{M_{XY}}{M_{Z}}-2\log\frac{M_{\Sigma}}{M_{Z}}+6\log\frac{M(\overline{3},3)}{\Lambda}$
$-5\log\frac{M(3,1)}{\Lambda}-6\log\frac{M(3,2)}{\Lambda}+\log\frac{M(6,1)}{\Lambda}$
$+4\log\frac{M(8,2)}{\Lambda}\biggr{]}.$ (11)
@@
Later, in Fig. 5, we plot the allowed region where GCU is achieved and the proton lifetime bounds are satisfied on the ${\overline{M}}$-$M(\overline{3},1)$ plane in the scenario involving three light components.
@@ 4 Proton Decay
×$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \times\left(\frac{v}{v_{45}}\right)^{4},$ (22)
@@
In Fig. 5, we plot the allowed region where GCU is achieved and the proton lifetime bounds are satisfied, in the scenario with three light components, on the ${\overline{M}}$-$M(\overline{3},1)$ plane. The colored region is allowed. The dashed lines correspond to the bounds on $M(\overline{3},1)$ obtained from various proton decay channels in Table 4. The left side of the colored region is excluded by the constraint from proton decay via the $XY$ gauge boson exchange, while the right side is excluded by the fact that $M(8,2)$ is smaller than the current bound set by the LHC. 111 We exclude the area above the horizontal line of $M(\overline{3},1)=2.4\times 10^{18}$ GeV because the $(\overline{3},1)_{1/3}$ boson mass exceeds the Planck scale.
Fig. 5: Allowed region on the ${\overline{M}}$-$M(\overline{3},1)$ plane where GCU is realized and the proton lifetime bounds are satisfied, in the scenario where three components of the 45 Higgs boson are light. The colored region is allowed. The dashed lines correspond to the bounds on $M(\overline{3},1)$ obtained from various proton decay channels in Table 4, of which the $p\rightarrow\bar{\nu}K^{+}$ mode gives the most stringent bound of $M(\overline{3},1)>3.4\times 10^{13}$ GeV. The left side of the colored region is excluded by the constraint from proton decay via the $XY$ gauge boson exchange, while the right side is excluded by the fact that $M(8,2)$ is smaller than the current bound set by the LHC.
@@ 5 Conclusion
In the case where two components of the 45 Higgs boson are lighter than the other GUT particles, we have observed that GCU is achieved only when the $(\overline{3},3)$ and $(8,2)$ components are light, owing to their substantial contributions to the beta functions. We have showed the solutions to the GCU conditions as a curve on the $M(\overline{3},3)-M(8,2)$, ${\overline{M}}-M(8,2)$, ${\overline{M}}-M(\overline{3},3)$ planes.
In the case where three components of the 45 Higgs boson are lighter, we have found that the components $(\overline{3},3)$ and $M(8,2)$ tend to be light, similar to the two-light-components scenario. We have focused on the case where the components $(\overline{3},1)$, $(\overline{3},3)$, $(8,2)$ are light, because $(\overline{3},1)$ induces proton decay and so we can constrain the model from the proton lifetime bounds. We have estimated proton lifetimes under the assumption that the 45 Higgs boson couples exclusively to the second generation fermions so that the fermion mass matrices show the pattern of the Georgi-Jarlskog model. We have revealed that $p\rightarrow\mu^{+}K^{0}$ is the dominant proton decay channel and that the Super-Kamiokande experiment constrains the mass of the $(\overline{3},1)$ component as $M(\overline{3},1)>3.4\times 10^{13}$ GeV. We have identified the viable parameter region on the ${\overline{M}}$-$M(\overline{3},1)$ plane where GCU is achieved and the proton lifetime bounds are satisfied.

### v3

@@ 2 Model: $SU(5)$ GUT with a 45 Higgs Field
$\mathcal{L}_{Y}=Y_{\bar{5}}\overline{\psi}^{a}\chi_{ab}\overline{5}_{H}^{b}+\frac{1}{8}Y_{10}\varepsilon_{abcde}\chi^{ab}\chi^{cd}5_{H}^{e}+Y_{\overline{45}}\overline{\psi}^{a}\chi_{ab}\overline{45}_{H}^{b}+\text{h.c.}$ (3)
@@ 3 Gauge Coupling Unification
$Q_{i}=\begin{pmatrix}u_{i}\\ d_{i}\end{pmatrix}_{L}$ $(3,2)_{1/6}$ 22 33 $1/5$
$u_{i}^{~c}=(u_{i}^{~c})_{L}$ $(\overline{3},1)_{-2/3}$ 11 0 $8/5$
$d_{i}^{~c}=(d_{i}^{~c})_{L}$ $(\overline{3},1)_{1/3}$ 11 0 $2/5$
$L_{i}=\begin{pmatrix}\nu_{ei}\\ e_{i}\end{pmatrix}_{L}$ $(1,2)_{-1/2}$ 0 11 $3/5$
$e_{i}^{~c}=(e_{i}^{~c})_{L}$ $(1,1)_{1}$ 0 0 $6/5$
$H=\begin{pmatrix}\phi_{+}\\ \phi_{0}\end{pmatrix}$ $(1,2)_{-1/2}$ 0 $1/6$ $1/10$
$H^{c}=\begin{pmatrix}H_{1}\\ H_{2}\\ H_{3}\end{pmatrix}$ $(\overline{3},1)_{1/3}$ $1/6$ 0 $1/15$
$(X^{\mu},Y^{\mu})$ $(3,2)_{-5/6}$ 7-7 21/2-21/2 35/2-35/2
$\Sigma_{8}$ $(8,1)_{0}$ $1/2$ 0 0
$\Sigma_{3}$ $(1,3)_{0}$ 0 $1/3$ 0
$\Sigma_{24}$ $(1,1)_{0}$ 0 0 0
$(1,2)_{-1/2}$ $(1,2)_{-1/2}$ 0 $1/6$ $1/10$
$(\overline{3},1)_{1/3}$ $(\overline{3},1)_{1/3}$ $1/6$ 0 $1/15$
$(\overline{3},3)_{1/3}$ $(\overline{3},3)_{1/3}$ $1/2$ 22 $1/5$
$(3,1)_{-4/3}$ $(3,1)_{-4/3}$ $1/6$ 0 $16/15$
@@
$\alpha_{1}^{-1}(M_{Z})=\alpha_{5}^{-1}(\Lambda)+\frac{1}{2\pi}\biggl[$ $\frac{41}{10}\log\frac{M_{Z}}{\Lambda}+\frac{1}{15}\log\frac{M_{H_{C}}}{\Lambda}-\frac{35}{2}\log\frac{M_{XY}}{\Lambda}+\frac{1}{10}\log\frac{M(1,2)}{\Lambda}$
@@
$\alpha_{2}^{-1}(M_{Z})-\frac{2}{12\pi}=\alpha_{5}^{-1}(\Lambda)+\frac{1}{2\pi}\biggl[$ $-\frac{19}{6}\log\frac{M_{Z}}{\Lambda}-\frac{21}{2}\log\frac{M_{XY}}{\Lambda}+\frac{1}{3}\log\frac{M_{\Sigma}}{\Lambda}+\frac{1}{6}\log\frac{M(1,2)}{\Lambda}$
@@
$\alpha_{3}^{-1}(M_{Z})-\frac{3}{12\pi}=\alpha_{5}^{-1}(\Lambda)+\frac{1}{2\pi}\biggl[$ $-7\log\frac{M_{Z}}{\Lambda}+\frac{1}{6}\log\frac{M_{H_{C}}}{\Lambda}-7\log\frac{M_{XY}}{\Lambda}+\frac{1}{2}\log\frac{M_{\Sigma}}{\Lambda}$
@@
[…]
@@
$-5\alpha_{1}^{-1}(M_{Z})+3\alpha_{2}^{-1}(M_{Z})+2\alpha_{3}^{-1}(M_{Z})=-\frac{1}{2\pi}\biggl[$ $42\log\frac{M_{XY}}{M_{Z}}+2\log\frac{M_{\Sigma}}{M_{Z}}+6\log\frac{M(\overline{3},3)}{\Lambda}-5\log\frac{M(3,1)}{\Lambda}$
$-6\log\frac{M(3,2)}{\Lambda}+\log\frac{M(6,1)}{\Lambda}+4\log\frac{M(8,2)}{\Lambda}\biggr].$ (11)
@@
Later, in Fig. 5, we plot the allowed region where GCU is achieved and the proton lifetime bounds are satisfied on the $\overline{M}$$M(\overline{3},1)$ plane in the scenario involving three light components.
@@ 4 Proton Decay
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \times\left(\frac{v}{v_{45}}\right)^{4},$ (22)
@@
In Fig. 5, we plot the allowed region where GCU is achieved and the proton lifetime bounds are satisfied, in the scenario with three light components, on the $\overline{M}$$M(\overline{3},1)$ plane. The colored region is allowed. The dashed lines correspond to the bounds on $M(\overline{3},1)$ obtained from various proton decay channels in Table 4. The left side of the colored region is excluded by the constraint from proton decay via the $XY$ gauge boson exchange, while the right side is excluded by the fact that $M(8,2)$ is smaller than the current bound set by the LHC.  [^1] We exclude the area above the horizontal line of $M(\overline{3},1)=2.4\times 10^{18}$ GeV because the $(\overline{3},1)_{1/3}$ boson mass exceeds the Planck scale.
Fig. 5: Allowed region on the $\overline{M}$$M(\overline{3},1)$ plane where GCU is realized and the proton lifetime bounds are satisfied, in the scenario where three components of the 45 Higgs boson are light. The colored region is allowed. The dashed lines correspond to the bounds on $M(\overline{3},1)$ obtained from various proton decay channels in Table 4, of which the $p\rightarrow\bar{\nu}K^{+}$ mode gives the most stringent bound of $M(\overline{3},1)>3.4\times 10^{13}$ GeV. The left side of the colored region is excluded by the constraint from proton decay via the $XY$ gauge boson exchange, while the right side is excluded by the fact that $M(8,2)$ is smaller than the current bound set by the LHC.
@@ 5 Conclusion
[…]
In the case where three components of the 45 Higgs boson are lighter, we have found that the components $(\overline{3},3)$ and $M(8,2)$ tend to be light, similar to the two-light-components scenario. We have focused on the case where the components $(\overline{3},1)$, $(\overline{3},3)$, $(8,2)$ are light, because $(\overline{3},1)$ induces proton decay and so we can constrain the model from the proton lifetime bounds. We have estimated proton lifetimes under the assumption that the 45 Higgs boson couples exclusively to the second generation fermions so that the fermion mass matrices show the pattern of the Georgi-Jarlskog model. We have revealed that $p\rightarrow\mu^{+}K^{0}$ is the dominant proton decay channel and that the Super-Kamiokande experiment constrains the mass of the $(\overline{3},1)$ component as $M(\overline{3},1)>3.4\times 10^{13}$ GeV. We have identified the viable parameter region on the $\overline{M}$$M(\overline{3},1)$ plane where GCU is achieved and the proton lifetime bounds are satisfied.